# Selective Variable Convolution Meets Dynamic Content Guided Attention for Infrared Small Target Detection

### v1

@@ IV-C Visualization Analysis
To qualitatively validate the advantages of DCGANet, we analyze representative samples from the three datasets (Fig. 7 - 8). Traditional model-driven methods, which rely on handcrafted features, exhibit high false alarms rates (Fig. 7, yellow boxes) and frequent missed detections (Fig. 8, cyan boxes). In contrast, deep learning-based methods leverage learned hierarchical representations to distinguish targets from complex background clutter effectively, showcasing superior detection robustness.
Gradient-weighted class activation mapping (GradCAM) visualizations for DCGANet (Fig. 9) highlight the focus for each module during the target extraction. The Selective Variable Convolution (SVC) employs diverse convolution forms compared to the baseline UNet, enabling robust feature preservation and edge refinement. The dynamic content guided attention (DCGA) module adaptively refines the distinction between target and background, demonstrating enhanced precision in focusing on true targets compared to static attention mechanisms. Additionally, the Adaptive Dynamic Feature Fusion (ADFF) module improves edge preservation, achieving sub-pixel accuracy in localization. These results collectively demonstrate that DCGANet’s modular, with each component exhibiting clear and discriminative attention patterns.
Figure 7: Visualization comparison of detection results via different methods on representative images from NUAA-SIRST, IRSTD-1K, and NUDT-SIRST datasets, indicate the land forests, oceans skys, and city building interfere. The red, green, and yellow boxes denote correct detection, false alarms, and missed detections, respectively.
Figure 8: Visualization comparesion the detection results of different methods (The red, green, and yellow boxes denote correct detection, false alarms, and missed detections). 3D mesh plot highlighting the features of small targets in the image.
@@ IV-D Ablation Study
IV-D1 Model-wise performance
Ablation experiments were conducted to evaluate the proposed SVC, DCGA, and ADFF modules, demostrating their individual and integrated impacts on model performance, as shown in Tab. IV. The baseline U-Net was used for comparison analysis.
@@ IV-D1 Model-wise performance
    Independent efficacy: Each module yields distinct performance improvements, corroborating its theoretical motivations (e.g. SVC’s 5.50% IoU gain on IRSTD-1K and DCGA’s 7.83% improvement).
@@
    Convolutional Advancement: Replacing vanilla convolutions with the SVC module elevates IoU by 4.62% on NUAA-SIRST and 5.50% on IRSTD-1K, demonstrating its superiority feature encoding multi-scale target features and suppressing background clutter.
@@
    Attention Mechanism Superiority:The DCGA outperforms established attention modules (e.g. SE, DAM, and CBAM) by adaptively modeling channel-specific spatial importance maps (SIMs) critical for sub-pixel infrared targets.
  4. 4.
    Contextual Fusion:The ADFF module’s dynamic feature fusion strategy, enhances detection accuracy, achieving a 9.62% improvement over baseline fusion strategies on the IRSTD-1K dataset. This demonstrates the effectiveness of leveraging broader contextual information for target detection in complex infrared scenes.
@@
Furthermore, Fig. 9 presents a qualitative comparison of detection results under different model configurations. The visualizations include original infrared inputs, ground-truth masks, attention maps from the baseline U-Net, and results from the progressively integrated DCGANet, showcasing the incremental benefits of each module in target localization and feature enhancement.
Figure 9: Qualitative comparison of detection results on infrared images. From left to right: original images, predictions by U-Net, U-Net+SVC, U-Net+SVC+DCGA, U-Net+SVC+DCGA+ADFF/DCGANet, and ground-truth masks. The improvements in target localization and detail are evident with the addition of each module.
@@
TABLE IV: Quantitative Ablation Study on DCGANet (selective variable convolution is marked SVC, dynamic content guided attention is marked DCGA, adaptive dynamic feature fusion module is marked ADFF), The $\uparrow$ indicates that higher values are better.
Strategy Module NUAA-SIRST IRSTD-1K NUDT-SIRST
SVC DCGA ADFF IoU $\uparrow$ nIoU $\uparrow$ IoU $\uparrow$ nIoU $\uparrow$ IoU $\uparrow$ nIoU $\uparrow$
  (a) 75.12 75.35 64.42 62.54 89.81 89.58
(b) 79.74 78.53 71.92 69.51 92.42 92.39
(c) 80.11 78.75 72.25 70.21 93.02 92.88
(d) 81.34 79.01 74.86 72.42 93.98 93.52
(e) $\checkmark$ $\checkmark$ $\checkmark$ 82.03 79.81 75.59 72.16 94.98 94.46
IV-D2 Ablation of SVC module
The SVC module integrates deformable convolutio (DC) and multi-dilation convolution (MDC) alongside Standard Convolution to generate more robust feature representations. Ablation results in Tab. V reveal that while DC alone provides a moderate performance gain, the integration of MDC yields significantly greater improvementds. This demonstrates the superior efficacy of the SVC composite convolutional block, where multi-rate dilated convolutions play a pivotal role in expanding receptive fields and capturing hierarchical contextual information.
These findings underscore the critical importance of broad receptive fields in IRSTD, serving as a key motivator for the design of the ADFF module. By leveraging broader spatial context, the ADFF module is engineered to refine target detection further, enhancing both accuracy and robustness in complex infrared scenes. This synergy between the SVC’s multi-convolutional feature extraction and the ADFF’s contextual fusion highlights the importance of hierarchical receptive field design in overcoming the challenges of low SNR and ambiguous target boundaries.
TABLE V: Ablation experiments of convolution branches of SVC.
Strategy Module NUAA-SIRST IRSTD1K NUDT-SIRST
SC MDC DC IoU $\uparrow$ nIoU $\uparrow$ IoU $\uparrow$ nIoU $\uparrow$ IoU $\uparrow$ nIoU $\uparrow$
(a) 80.35 77.59 72.14 70.12 91.21 91.96
(b) 81.72 78.47 74.01 70.85 93.30 93.88
(c) 80.97 78.22 73.45 70.36 92.64 93.37
(d) 82.03 79.81 75.59 72.16 94.98 94.46
IV-D3 Ablation of DCGA module
Traditional attention modules (SE/DAM/CBAM) fail to address the challenges specific to IRSTD:
  1. 1.
    Under the condition of low SNR, single-channel simulation cannot adequately represent cross-channel variance.
  2. 2.
    Compared with the serial design, the channel shuffling architecture of parallel processing DCGA improves the small target recognition rate.
As shown in Tab. VI, in detection performance, DCGA has significantly improved compared with SE, DAM and CBAM. This shows that DCGA is able to better recalibrate the feature distribution by learning Spatial Importance Maps (SIM) to focus on signal differences between channel-wise, which is particularly critical for enhancing infrared small target features.
@@ IV-D3 Ablation of DCGA module
TABLE VI: Ablation of DCGA compared with other mainstream attention mechanisms
Strategy NUAA-SIRST IRSTD1K NUDT-SIRST
IoU $\uparrow$ nIoU $\uparrow$ IoU $\uparrow$ nIoU $\uparrow$ IoU $\uparrow$ nIoU $\uparrow$
SE 78.74 77.90 68.13 68.47 90.93 90.98
DAM 79.88 78.01 70.19 69.11 91.29 91.65
CBAM 80.00 77.88 71.20 68.59 92.30 92.04
DCGA 82.03 79.81 75.59 72.16 94.98 94.46
TABLE VII: Ablation study of DCGA. All the experiments are conducted on three datasets.
Method DCGAcas DCGAw/o r DCGAw/o CS DCGA
parallel CA&SA
cascaded CA&SA
refine ✓(remove CS)
NUAA-SIRST IoU 80.12 75.15 78.54 82.03
nIoU 78.71 75.24 77.12 79.81
IRSTD-1K IoU 72.25 64.41 71.34 75.59
nIoU 70.20 62.97 69.04 72.16
NUDT-SIRST IoU 93.04 89.14 91.19 94.48
nIoU 92.78 89.43 91.01 94.46
Further ablation studies on the DCGA design settings are as follows:
  1. 1.
    DCGAcas represents the cascaded connection of channel attention (CA) and spatial attention (SA) in contrast to the parallel mode of DCGA;
  2. 2.
    DCGAw/o r denotes the parallel connection of channel attention (CA) and spatial attention (SA) but without the refinement process;
  3. 3.
    DCGAw/o CS indicates the parallel connection of channel attention (CA) and spatial attention (SA) but excludes the Channel Shuffle (CS) operation in the refinement stage;
  4. 4.
    DCGA is the proposed method that integrates parallel connection of channel attention (CA) and spatial attention (SA), along with a complete refinement process that includes channel shuffle (CS).
The results are presented in Tab. VII, which reveals that model performance slightly declines when the channel shuffle (CS) operation is removed. This result indicates that CS plays a crucial role in facilitating cross-channel information interaction to highlight small target features. Additionally, the channel attention (CA) and spatial attention (SA) mechanisms in DCGA are deployed in parallel, efficiently modeling the salient regions of small targets by generating initial spatial importance maps (SIMs) in a parallel manner. We also experimented with a cascaded design, and the results show that both parallel and cascaded designs exhibit very similar performance in terms of signal-to-noise ratio (SNR) gain. However, when the refinement process is omitted (DCGAw/o CS), performance significantly deteriorates, further confirming the importance of the refinement stage in enhancing the feature representation of infrared small targets.
@@ V Conclusion
In this paper, we propose a dynamic content guided attention multiscale feature aggregation network (DCGANet), which adopts a “coarse-to-fine” attention principle to achieve high detection accuracy. First, the selective variable convolution (SVC) module is introduced, integrating standard convolution, deformable convolution, and multi-rate dilated convolution. This module expands the receptive field and enhances non-local features, thereby effectively improving the discrimination between targets and backgrounds. Second, the core component of DCGANet is the dynamic content guided attention (DCGA) module, which uses two-stage attention to first direct the network’s focus to significant regions in feature maps and then determine whether these regions correspond to targets or background interference. By retaining the most salient responses, this mechanism suppresses false alarms effectively. Additionally, the adaptive dynamic feature fusion (ADFF) module is proposed to substitute for static feature cascading. The dynamic feature fusion strategy enables DCGANet to adaptively integrate contextual features, enhancing its capability to discriminate between true targets and false alarms. DCGANet establishes new benchmarks across multiple datasets.

### v2

@@ IV-C Visualization Analysis
To qualitatively validate the advantages of DCGANet, we analyze representative samples from the three datasets (Figures 7 and 8). Traditional model-driven methods, which rely on handcrafted features, exhibit high false alarm rates (yellow boxes) and frequent missed detections. In contrast, deep learning-based methods leverage learned hierarchical representations to distinguish targets from complex background clutter more effectively, showcasing superior detection robustness.
Gradient-weighted Class Activation Mapping (GradCAM) visualizations for DCGANet (Figure 9) highlight the focus area for each module during target extraction. The SVC module employs diverse convolutional forms, enabling robust feature preservation and edge refinement compared to the baseline. The DCGA module adaptively refines the distinction between target and background, demonstrating enhanced precision in focusing on true targets. Finally, the ADFF module improves edge preservation by dynamically fusing features, achieving sub-pixel accuracy in localization. These results collectively demonstrate that DCGANet’s modules operate synergistically, with each component exhibiting clear and discriminative attention patterns.
Figure 7: Visualization comparison of detection results on representative images from the three datasets, showcasing interference from land, oceans, and cityscapes. The red, green, and yellow boxes denote correct detections, false alarms, and missed detections, respectively.
Figure 8: Further visualization comparisons. The main images show detection results, while the 3D mesh plots below each one highlight the intensity profile of the small target feature against its local background.
Figure 9: Qualitative comparison of attention maps on infrared images. From left to right: original images, predictions by U-Net (baseline), U-Net+SVC, U-Net+SVC+DCGA, U-Net+SVC+DCGA+ADFF (our full DCGANet), and ground-truth masks. The improvements in target localization and detail suppression of noise are evident with the addition of each module.
@@ IV-D Ablation Study
IV-D1 Component-wise Performance Analysis
Ablation experiments were conducted to systematically evaluate the proposed SVC, DCGA, and ADFF modules, demonstrating their individual and integrated impacts on model performance. As shown in Table IV, the baseline U-Net was used for comparison.
@@ IV-D1 Component-wise Performance Analysis
    Independent Efficacy: Each module individually yields distinct performance improvements over the baseline (a), corroborating its theoretical motivation. For instance, adding DCGA alone (c) improves IoU by 7.67% on IRSTD-1K.
@@
    Effectiveness of SVC: Replacing standard convolutions with our SVC module (b) elevates the IoU by 4.53% on NUAA-SIRST and 7.35% on IRSTD-1K, proving its superiority in encoding multi-scale target features while suppressing clutter.
@@
    Synergy of Modules: Combining SVC and DCGA (d) provides a significant boost over using either module alone. The full model, DCGANet (e), which also includes ADFF, achieves the overall best performance, demonstrating the powerful synergy of the three proposed components.
@@
Furthermore, Figure 9 presents a qualitative comparison of attention maps under different model configurations. The visualizations showcase the incremental benefits of each module in refining target localization and enhancing feature representation.
@@
TABLE IV: Quantitative ablation study of DCGANet’s core components. We measure the performance improvement by sequentially adding Selective Variable Convolution (SVC), Dynamic Context-Guided Attention (DCGA), and Adaptive Dynamic Feature Fusion (ADFF) to a U-Net baseline.
[…]
(b) 76.59 75.42 69.07 66.76 88.76 88.73
(c) 76.94 75.64 69.39 67.43 89.34 89.20
(d) 78.12 75.88 71.89 69.55 90.26 89.82
(e) 78.78 76.65 72.60 69.31 91.22 90.72
[…]
IV-D2 Ablation of SVC Module Design
The SVC module integrates Standard Convolution (SConv), Deformable Convolution (DConv), and Multiple Dilation Convolution (MDConv). To analyze each branch’s contribution, we conduct an ablation study shown in Table V. The results reveal that while adding DConv provides a moderate performance gain (c), the addition of MDConv yields significantly greater improvements (b). The full three-branch SVC (d) achieves the best performance, demonstrating that multi-rate dilated convolutions play a pivotal role in expanding the receptive field and capturing crucial hierarchical context.
TABLE V: Ablation study of the convolutional branches within the SVC module. We start with a standard convolution (SConv) and add the other branches. Results on NUDT-SIRST are included for comprehensive analysis.
[…]
(b) 78.49 75.36 71.08 68.04 90.46 90.31
(c) 77.76 75.13 70.54 67.57 89.25 89.03
(d) 78.78 76.65 72.60 69.31 91.22 90.72
[…]
IV-D3 Ablation of DCGA Module Design
We compare DCGA against other mainstream attention mechanisms (SE, DAM, CBAM) in Table VI. DCGA significantly outperforms them, showing its ability to better recalibrate feature distributions by learning channel-specific Spatial Importance Maps (SIMs), which is critical for enhancing infrared small target features.
@@ IV-D3 Ablation of DCGA Module Design
TABLE VI: Comparison of DCGA with other mainstream attention mechanisms.
[…]
DAM 76.71 74.92 67.41 66.38 87.67 88.02
CBAM 76.83 74.79 68.38 65.88 88.64 88.40
DCGA (Ours) 78.78 76.65 72.60 69.31 91.22 90.72
[…]
Furthermore, we conducted a detailed ablation of the DCGA’s internal design choices (Table VII). The study validates our design:
  • DCGAcas: A cascaded connection of Channel (CA) and Spatial Attention (SA) performs slightly worse than our parallel design, suggesting parallel processing is more effective for this task.
  • DCGAw/o r: Removing the refinement stage entirely causes a significant performance drop, confirming the importance of the refinement process.
  • DCGAw/o CS: Removing only the Channel Shuffle (CS) from the refinement stage also degrades performance, indicating that CS is crucial for cross-channel information interaction.
TABLE VII: Detailed ablation study on the internal design of the DCGA module. We compare our full design against three variants: a cascaded version (DCGAcas), one without the refinement stage (DCGAw/o r), and one without the channel shuffle operation (DCGAw/o CS). The best results are highlighted in bold.
[…]
nIoU (%) $\uparrow$ 75.60 72.26 74.07 76.65
[…]
nIoU (%) $\uparrow$ 67.42 60.48 66.31 69.31
[…]
nIoU (%) $\uparrow$ 89.10 85.89 87.41 90.72
[…]
@@ V Conclusion
In this paper, we propose a dynamic content-guided attention multiscale feature aggregation network (DCGANet), which adopts a "coarse-to-fine" attention principle to achieve high detection accuracy. First, the selective variable convolution (SVC) module is introduced, integrating standard convolution, deformable convolution, and multi-rate dilated convolution. This module expands the receptive field and enhances non-local features, thereby effectively improving the discrimination between targets and backgrounds. Second, the core component of DCGANet is the Dynamic Content-Guided Attention (DCGA) module, which uses two-stage attention to first direct the network’s focus to significant regions in feature maps and then determine whether these regions correspond to targets or background interference. By retaining the most salient responses, this mechanism suppresses false alarms effectively. Additionally, the Adaptive Dynamic Feature Fusion (ADFF) module is proposed to substitute for static feature cascading. The dynamic feature fusion strategy enables DCGANet to adaptively integrate contextual features, enhancing its capability to discriminate between true targets and false alarms. DCGANet establishes new benchmarks across multiple datasets.